OpenStudy (anonymous):

In a function, how exactly can you figure out the y intercept without graphing?

OpenStudy (abb0t):

example?

OpenStudy (4n1m0s1ty):

Just find the value where x is equal to zero for a function f(x)

OpenStudy (anonymous):

y = 7/(x-9) + 2

OpenStudy (anonymous):

PARTY

hartnn (hartnn):

to get y intercept, just plug in x=0 in the function

OpenStudy (cwrw238):

fnd the value of when x = 0

OpenStudy (cwrw238):

* value of y

hartnn (hartnn):

so what u get when u put x=0 in y = 7/(x-9) + 2 ?

OpenStudy (anonymous):

y = 2?

hartnn (hartnn):

ummm..no, try again ? y = 7/(0-9)+2 = ... ?

OpenStudy (anonymous):

Ohhh, 11/9 xD

hartnn (hartnn):

yup :) thats your y-intercept (0,11/9) :)

OpenStudy (anonymous):

Oh :o